This problem is from a test. I have the solution, but I am unsure of where one of the numbers came from, so I would not be able to do a similiar problem on the final.
The class average on a math test was 82 out 100. The girls in the class scored 966 points. How many total points did the 8 boys score?
The solution I have looks like this:
966 + b/19 = 82
966 + b = 1558 (where does the 1558 come from?)
b = 592
: This problem is from a test. I have the solution, but I am unsure of where one of the numbers came from, so I would not be able to do a similiar problem on the final.
The class average on a math test was 82 out 100. The girls in the class scored 966 points. How many total points did the 8 boys score?
The solution I have looks like this:
966 + b/19 = 82
966 + b = 1558 (where does the 1558 come from?)
b = 592

Answer by vleith(1238)

Your problem did not state how many students were in the class. But your first line in the solution uses the number 19. So I will assume there are 19 students.
You are told the class average is 82, that girls total was 966 and that 8 boys took the test.
So, the total points scored on the test is:
Total= girlsTotal + boysTotal
Total = 966 + 8*boysAverage
You also know the class average was 82. So the Total= 19*82 = 1558
Substitute the total and solve for boysAverage
Total = 966 + 8*boysAverage
1558 = 966 + 8*boysAverage
1558-966 = 8*boysAverage
592 = 8*boysAverage
74 = boysAverage